# FusionPortableV2: A Unified Multi-Sensor Dataset for Generalized SLAM Across Diverse Platforms and Scalable Environments

### v1

@@ Abstract
Simultaneous Localization and Mapping (SLAM) technology has been widely applied in various robotic scenarios, from rescue operations to autonomous driving. However, the generalization of SLAM algorithms remains a significant challenge, as current datasets often lack scalability in terms of platforms and environments. To address this limitation, we present FusionPortableV2, a multi-sensor SLAM dataset featuring notable sensor diversity, varied motion patterns, and a wide range of environmental scenarios. Our dataset comprises $27$ sequences, spanning over $2.5$ hours and collected from four distinct platforms: a handheld suite, wheeled and legged robots, and vehicles. These sequences cover diverse settings, including buildings, campuses, and urban areas, with a total length of $38.7km$. Additionally, the dataset includes ground-truth (GT) trajectories and RGB point cloud maps covering approximately $0.3km^{2}$. To validate the utility of our dataset in advancing SLAM research, we assess several state-of-the-art (SOTA) SLAM algorithms. Furthermore, we demonstrate the dataset’s broad applicability beyond traditional SLAM tasks by investigating its potential for monocular depth estimation. The complete dataset, including sensor data, GT, and calibration details, is accessible at https://fusionportable.github.io/dataset/fusionportable_v2.
@@
Dataset, SLAM, General Robotic, LiDAR, Camera, Kinematic, IMU, Encoder
@@ 1.1 Motivation
Real-world robotics datasets play a crucial role in the development of SLAM. They not only cover a wide range of environments but also feature intricately designed sequences that pose challenges. They serve as essential resources for thoroughly training and evaluating SLAM systems and circumventing the reliance on complex and potentially imprecise physical simulations. Moreover, such datasets eliminate the need for budget and workforce, such as system integration, calibration, and extensive field operations (Nguyen et al., 2022). These accessible resources allow for broader participation in SLAM research, facilitating development of novel algorithms without relying on high costs or specialized expertise. Currently, datasets become particularly important since the field of robotics research is experiencing a paradigm shift from traditional handcrafted and model-based methodologies to data-driven approaches (Brohan et al., 2022; Shah et al., 2023b), alongside the emergence of hybrid strategies.
The robustness and performance of large-scale and long-term SLAM strongly rely on a robot’s perceptual capability, especially in field environments (e.g., subterranean (Reinke et al., 2022) and forest (Knights et al., 2023)). As outlined in Table 1, SLAM datasets, encompassing a wide range of scenarios from both simulated and real-world environments, are increasingly being augmented with multi-modal sensor data, moving beyond the traditional reliance on singular modal inputs like images. For example, cameras capture dense and high-resolution 2D images containing texture and pattern information of surroundings. However, cameras are vulnerable to adverse illumination conditions (e.g., darkness and glare) due to their passive nature in measuring. In contrast, range sensors such as LiDARs and Radars provide sparse but highly accurate structural information by exploiting their respective light sources. Integrating cameras with range sensors often yields more reliable results across a variety of perception tasks compared to relying on a single sensor type (Lai et al., 2022). Therefore, the complementary strengthes offered by various sensors drives the exploration of novel sensor fusion algorithms (Lin and Zhang, 2022), enabling robots to autonomously navigate through diverse environments with increased accuracy and robustness.
Despite the advancements in SLAM, there remains a notable deficiency in the diversity and breadth of available SLAM datasets, especially concerning variations in environments, sensor modalities, and platforms executing diverse motions. This limitation affects the algorithms’ flexibility and applicability in real-world scenarios, underlining an immediate need for more inclusive datasets. Drawing on the recent success in training generalized manipulation (Brohan et al., 2022) and navigation (Shah et al., 2023a) models, it is our conviction that datasets featuring high motion and environmental diversity are crucial for the development of a versatile, generalized SLAM system.
@@ 1.2 Contributions
Following this route, this paper aims to introduce a comprehensive multi-sensor dataset with a standardized format, share insights and methodologies from our data collection process, and provide benchmarking tools along with GT for evaluating various algorithms. Building upon our previous open-source FusionPortable dataset (Jiao et al., 2022), we present FusionPortableV2, an extensive upgrade that broadens the dataset in terms of data modalities, scenarios, and motion capabilities. The key contributions of this paper are threefold:
@@
    We introduce significant improvements to the FusionPortable dataset by extending the spectrum of platforms to encompass high-speed vehicles and spanning more than $12$ types of environments (e.g., campuses, underground areas, parking lots, highways, etc.). These enhancements not only enrich the dataset’s diversity and complexity but also bring substantial upgrades in the precision of ground truth data and the integration of raw kinematic data collected from ground robots. Our collection encompasses $27$ sequences, spanning $2.5$ hours and covering a total distance of $38.7km$.
@@
    We furnish detailed ground truth trajectories and maps, encompassing an area of around $0.3km^{2}$, for most of the sequences. This facilitates a more thorough algorithm evaluation and supports a wide range of research objectives. The rich variety of platforms and settings in our dataset exemplifies its utility as a benchmark for assessing various navigation algorithms beyond SLAM.
  3. 3.
    We have expanded the utility of the dataset by incorporating experiments on monocular depth estimation, addressing both the training and evaluation stages, in addition to localization and mapping. Throughout the process of platform development and data collection, we have tackled numerous technical challenges and have meticulously described the encountered issues and their solutions. This guidance is intended to serve as a valuable resource for future researchers in the field. To foster collaborative advancements, we have publicly released all data and implementation details, aiming to open up extensive research avenues in field robotics and contribute to the development of versatile and resilient robotic systems.
@@ 1.3 Organization
The remainder of this paper is structured in the following manner: Section 2 discusses related works most of on SLAM datasets and summarizes key contributions of this paper. Section 3 outlines the hardware setup and sensor details. Section 4 covers sensor calibration procedures. Section 5 describes the dataset, including platform characteristics and scenarios. Section 6 introduces details post-processing steps on raw sensor measurements and GT data. Section 7 presents the methodologies used for evaluating localization, mapping, and mocular depth estimation. Known issues of this dataset are also dicussed. Finally, Section 8 concludes the paper and suggests directions for future research.
@@ 2 Related Works
In the last decade, high-quality datasets have crucially driven SLAM algorithm advancements, yielding significant time and cost savings in platform development and data acquisition. The rapid progress in sensor and robotics technology has led to the widespread adoption of multi-modal sensors across various robotic platforms, as detailed in Table 1. This evolution has set new benchmarks and hastened the enhancement of SOTA algorithms, covering both handcrafted and data-driven methods like VINS-Mono (Qin et al., 2018), FAST-LIO2 (Xu et al., 2022), VILENS (Wisth et al., 2022), DROID-SLAM (Teed and Deng, 2021), and Gaussian Splatting SLAM (Matsuki et al., 2023).
2.1 Specific Platform Datasets
@@
A wide array of SLAM datasets predominantly highlights visual-inertial fusion, targeting specific mobile platforms and venues. Platfors are ranged from handheld devices (Pfrommer et al., 2017; Schubert et al., 2018; Zuñiga-Noël et al., 2020) to drones (Burri et al., 2016; Majdik et al., 2017), unmanned ground vehicles (Pire et al., 2019), and aquatic vehicles such as canoes (Miller et al., 2018). Among them, the UZH-FPV dataset (Delmerico et al., 2019) is featured by its extended, rapid trajectories from aggressive drone flight with the integration of event cameras.
@@
Urban environment datasets introduce specific challenges like adverse lighting, weather conditions, and larger scales. Long-range sensors, notably LiDARs and Radars, are preferred for their capabilities. The KITTI dataset (Geiger et al., 2013) sets a benchmark in autonomous driving with its rich urban sensor data collection, including LiDAR and cameras. Further developments in driving datasets have expanded across dimensions of duration (Maddern et al., 2017), urban complexity (Jeong et al., 2019), and weather adversity (Agarwal et al., 2020). The DSEC dataset (Gehrig et al., 2021), akin to UZH-FPV, leverages stereo event cameras for extensive driving scenes. Moreover, Radars are essential for outdoor perception, offering advantages in range, velocity measurement via the Doppler effect, and weather resilience. Radar-focused datasets like Boreas (Burnett et al., 2023), Oxford Radar Robocar (Barnes et al., 2020), and OORD (Gadd et al., 2024) collected data in conditions such as fog, rain, and snow.
@@
Campus-scene datasets such as NCLT (Carlevaris-Bianco et al., 2016), M2DGR (Yin et al., 2021), NTU-Viral (Nguyen et al., 2022), ALITA (Yin et al., 2022), and FusionPortable (Jiao et al., 2022) also pose challenges for SLAM, given their diverse environmental appearance and structure. A typical campus features a variety of environments, inculding dense vegetation, open spaces, and complex buildings with multiple levels and detailed layouts. The changing lighting, seasonal foliage variations, and movement of pedestrians and vehicles add complexity to campus environments. The NCLT dataset highlights these factors as crucial for life-long SLAM challenges. However, despite their value, these datasets, collected via specific platforms like unmanned ground vehicles (UGVs) and drones, fall short in showcasing diverse motion patterns, especially aggressive maneuvers.
The Newer College dataset (Ramezani et al., 2020), Hilti-Oxford Dataset (Zhang et al., 2022), and MARS-LVIG dataset (Li et al., 2024) demonstrated the portable multi-sensor suites for collecting multi-modal data in a wide range of environments. Driven by the evaluation need for reconstruction and $6$-DoF motion estimates, the former two datasets additionally provide dense 3D global maps and allow the evaluation of mapping performance.
2.2 Multiple Platforms Datasets
@@
Recent SLAM research trends emphasize algorithm generalization across different platforms and scales, aiming for motion characteristic integration from varied platforms with minimal parameter tuning for diverse scenarios. The MVSEC dataset (Zhu et al., 2018) pioneered in capturing multi-sensor data across platforms, excluding UGV sequences. Conversely, the Nebula dataset (Reinke et al., 2022), developed during the DARPA Subterranean Challenge, includes field environments with wheeled and legged robots, offering precise maps and trajectories but missing urban data and having varied sensor setups across platforms. The M3ED dataset (Chaney et al., 2023), although closely aligned with our objectives, lacks indoor data and platform-specific kinematic measurements, underscoring the unique contribution of our dataset.
@@
Table 2: The sensors used in this dataset and their corresponding specifications
@@
2.3 Robot-Agnostic General Models
Recent progress in natural language processing and visual perception has spurred research into universal, robot-agnostic models for zero-shot transfer or task-specific fine-tuning, highlighting the need for expansive datasets. The Open X-Embodiment dataset (Padalkar et al., 2023) demonstrates this by presenting over $500$ skills across $22$ robot types, emphasizing the crucial role of large-scale datasets in robotic manipulation advancement. The DROID dataset (Khazatsky et al., 2024), collected by 13 institutions using identical hardware configurations to maximize portability and flexibility, further contributing to the advancement of robotic manipulation research. Likewise, the General Navigation Model (GNM) (Shah et al., 2023a) underlines the value of flexible, scalable navigation methods. Trained on diverse robot datasets, GNM’s unified model excels in various settings, even with sensor and actuator issues. Our objective aligns with that of GNM, with a specific focus on datasets: we aim to offer diverse, high-quality data towards the development and benchmark of more generalized and more robust SLAM algorithms.
@@ 3 System Overview
Figure 2: Data collection framework, highlighting the data flow and synchronization processes. The Red arrow indicate PPS signals for synchronization, green arrows show UTC time synchronization, and blue arrows represent sensor triggering signals, and black arrows depict the flow of raw data.
This section presents our developed multi-sensor suite, designed for easy integration with various mobile platforms through plug-and-play functionality. All sensors are securely mounted on a precision-engineered aluminum alloy frame, facilitating a unified installation. Additionally, we detail the devices employed for collecting GT trajectories and maps.
@@ 3.1 Suite Setup and Synchronization
The Multi-Sensor Suite (MSS) integrates exteroceptive and proprioceptive sensors, including a 3D Ouster LiDAR, stereo frame and event cameras, and IMUs, depicted in its CAD model in Fig. 1. To manage data collection efficiently, we use two PCs synchronized via the Network Time Protocol (NTP) server. The primary PC processes data from the frame cameras and IMU, while the auxiliary PC handles additional data types. Both PCs are equipped with a $1$TB SSD, $64$GB of DDR4 memory, and an Intel i7777 processor, running Ubuntu with a real-time kernel patch and employing the Robot Operating System (ROS) for data collection. Subsequent sections will elaborate on our data synchronization approach and the specific features of each sensor.

### v2

@@ Abstract
Simultaneous Localization and Mapping (SLAM) technology has been widely applied in various robotic scenarios, from rescue operations to autonomous driving. However, the generalization of SLAM algorithms remains a significant challenge, as current datasets often lack scalability in terms of platforms and environments. To address this limitation, we present FusionPortableV2, a multi-sensor SLAM dataset featuring sensor diversity, varied motion patterns, and a wide range of environmental scenarios. Our dataset comprises $27$ sequences, spanning over $2.5$ hours and collected from four distinct platforms: a handheld suite, a legged robot, a unmanned ground vehicle (UGV), and a vehicle. These sequences cover diverse settings, including buildings, campuses, and urban areas, with a total length of $38.7km$. Additionally, the dataset includes ground-truth (GT) trajectories and RGB point cloud maps covering approximately $0.3km^{2}$. To validate the utility of our dataset in advancing SLAM research, we assess several state-of-the-art (SOTA) SLAM algorithms. Furthermore, we demonstrate the dataset’s broad application beyond traditional SLAM tasks by investigating its potential for monocular depth estimation. The completae dataset, including sensor data, GT, and calibration details, is accessible at https://fusionportable.github.io/dataset/fusionportable_v2.
@@
SLAM Dataset, Sensor Fusion, Mapping, Mobile Robots
@@ 1.1 Motivation
Real-world robotic datasets are vital for algorithm development. They provide diverse environments and challenging real-world sequences for training and evaluation of systems. They reduce costs and workforce requirements, such as system integration, calibration, and field operations (Nguyen et al., 2022), promoting broader participation in robotic research and fostering novel algorithm development. As robotics research transitions from traditional handcrafted methods to data-driven and hybrid approaches (Brohan et al., 2022; Shah et al., 2023b), the importance of these datasets continues to grow. Building upon this trend, this paper contributes to the exploration of SLAM dataset’s potential by introducing a diverse multi-sensor dataset. Our goal aims to develop a dataset to address the generalization challenges in robotic perception and navigation across various environments and operational conditions.
As outlined in Table 1, recent SLAM datasets exhibit two key trends: (1) they encompass a variety of large-scale real-world environments (e.g., urban roads (Burnett et al., 2023), subterranean (Reinke et al., 2022), and forests (Knights et al., 2023)), and (2) they incorporate heterogeneous sensors to enhance perceptual awareness. For example, cameras capture dense and high-resolution 2D images containing texture and pattern information of surroundings. However, cameras are vulnerable to adverse illumination conditions (e.g., darkness and glare) due to their passive nature in measuring. In contrast, range sensors such as LiDARs and Radars provide sparse but accurate structural information by exploiting their respective light sources. Integrating cameras with range sensors often yields more reliable results across a variety of perception tasks compared to relying on a single sensor type. Therefore, the complementary strengthen offered by various sensors drives the exploration of novel sensor fusion algorithms (Lin and Zhang, 2022), enabling robots to autonomously navigate through diverse environments with increased accuracy and robustness.
Despite the advancements in SLAM, a gap persists between the diversity of real-world scenarios and available datasets, especially concerning variations in environments, sensor modalities, and platforms executing diverse motions. This disparity affects the further development and evaluation of SLAM algorithms, potentially limiting their generalization capability and robustness in diverse real-world environments. Drawing on the recent success in the generalized manipulation and navigation models such as RT-2 (Brohan et al., 2023) and GNM (Shah et al., 2023a), it is our conviction that datasets featuring high motion and environmental diversity are crucial for the development of a versatile and generalized SLAM system.
@@ 1.2 Contributions
This paper aims to provide diverse, high-quality data using a consistent hardware setup, facilitating the development and benchmarking of robust and generalized SLAM systems. To achieve this objective, we present a comprehensive multi-sensor dataset, along with a detailed description of our data collection methodologies and complete benchmarking tools for evaluation. Building upon our previous FusionPortable dataset (Jiao et al., 2022), we introduce FusionPortableV2, a significant upgrade that expands the dataset in terms of data modalities, scenarios, and motion capabilities. This paper presents two main contributions:
@@
    Diversity: We have improved the FusionPortable dataset by expanding the range of platforms to include high-speed vehicles and to cover over $12$ types of environments (e.g., campuses, underground areas, parking lots, highways, etc.). These enhancements not only increase the dataset’s diversity and complexity but also substantially improve the accuracy of ground truth data and the integration of raw kinematic data collected from ground robots. Our collection includes $27$ sequences, spanning $2.5$ hours and covering a total distance of $38.7km$. We also provide valuable ground-truth trajectories and maps encompassing a campus area of approximately $0.3km^{2}$ to serve as benchmarks for evaluating various navigation algorithms beyond SLAM.
@@
    Versatile: We demonstrate the dataset’s value for training models and evaluating algorithms through experiments on SLAM and monocular depth estimation. These experiments highlight the challenges and the need for cross-platform adaptability in diverse environments. With its rich variety of platforms and settings, the dataset can be extended to benchmarks for tasks such as cross-view and cross-modality localization (Shi et al., 2023), environmental mapping (Kerbl et al., 2023; Hu et al., 2024a), and navigation tasks with the high-quality map as simulation.
During the platform development and data collection process, we addressed numerous technical challenges and meticulously documented the issues encountered and their solutions. This guidance should be a valuable resource for future researchers in the field. To promote collaborative advancements, we have publicly released all data and implementation details. We believe this will open up a wide range of research opportunities in field robotics and aid in the development of versatile, resilient robotic systems.
@@ 1.3 Organization
The remainder of this paper is structured in the following manner: Section 2 discusses related works most of on SLAM datasets and summarizes key contributions of this paper. Section 3 outlines the hardware setup and sensor details. Section 4 covers sensor calibration procedures. Section 5 describes the dataset, including platform characteristics and scenarios. Section 6 introduces details post-processing steps on raw sensor measurements and GT data. Section 7 presents the methodologies used for evaluating localization, mapping, and monocular depth estimation. Known issues of this dataset are also discussed. Finally, Section 8 concludes the paper and suggests directions for future research.
@@ 2 Related Works
In the last decade, the availability of high-quality datasets has significantly accelerated the development of SOTA SLAM algorithms by reducing the time and cost associated with data acquisition and algorithm evaluation. The rapid progress in sensor and robotics technology has led to the widespread adoption of multiple sensors across various robotic platforms. This evolution has set new benchmarks and hastened the enhancement of SOTA algorithms, spanning both handcrafted and data-driven methods such as VINS-Mono (Qin et al., 2018), FAST-LIO2 (Xu et al., 2022), VILENS (Wisth et al., 2022), DROID-SLAM (Teed and Deng, 2021), and Gaussian Splatting SLAM (Matsuki et al., 2024).
Recent advancements in the SLAM field have also extended to areas such as place recognition and collaborative SLAM. Although these areas are not the primary focus of our work, they contribute significantly to the broader SLAM research community. Datasets such as Wild-Places (Knights et al., 2023) and HeLiPR (Jung et al., 2023) are specifically designed for LiDAR-based place recognition, emphasizing large-scale, long-term localization under varying appearance conditions (Yin et al., 2024). Collaborative SLAM, which enables information sharing and cooperative mapping among multiple robots, has gained attention with datasets such as Kimera-Multi (Tian et al., 2023), GrAco (Zhu et al., 2023), and S3E (Feng et al., 2022). These datasets provide multi-sensor data (e.g., cameras and LiDARs) and benchmarking tools, which are useful for both collaborative and single-robot SLAM evaluation. However, as they focus primarily on multi-robot scenarios with identical platforms, their utility for exploring cross-platform variability and single-robot SLAM challenges is limited.
In contrast, our FusionPortableV2 dataset is primarily designed for short-term odometry and SLAM, with a focus on accurately tracking the robot’s pose and constructing consistent maps within a single session or over a shorter time period. The dataset features diverse platforms, sensor configurations, and environments, making it better suited for studying SLAM generalization across different conditions compared to most related works, as detailed in Table 1.
2.1 Specific-Platform Datasets
@@
Early SLAM datasets predominantly focused on visual-inertial fusion, targeting specific platforms and environments. This focus was largely due to the ubiquity and convenience of visual and inertial sensors which are cheap and lightweight. They cover sequences which were captured by different platforms ranging from handheld devices (Pfrommer et al., 2017; Schubert et al., 2018; Zuñiga-Noël et al., 2020), drones (Burri et al., 2016; Majdik et al., 2017; Delmerico et al., 2019; Li et al., 2024), unmanned ground vehicles (Pire et al., 2019), and aquatic vehicles (Miller et al., 2018), respectively. Notably, the UZH-FPV dataset (Delmerico et al., 2019) stands out for its integration of event cameras and the inclusion of rapid trajectories from aggressive drone flights.
@@
Concurrently, in the automotive industry, urban environment datasets introduce specific challenges including adverse lighting, weather conditions, and larger scales. Long-range sensors such as LiDARs and Radars are preferred for their capabilities, even though they were initially bulky and costly. The KITTI dataset (Geiger et al., 2013) sets a benchmark in autonomous driving with its rich urban sensor data collection. Further developments in driving-related datasets have expanded across dimensions of duration (Maddern et al., 2017), urban complexity (Jeong et al., 2019), and weather adversity (Agarwal et al., 2020). The DSEC dataset (Gehrig et al., 2021), akin to UZH-FPV, leverages stereo event cameras for extensive driving scenes. Moreover, Radars are essential for outdoor perception, offering advantages in range, velocity measurement via the Doppler effect and weather resilience. Related datasets such as Boreas (Burnett et al., 2023), Oxford Radar RoboCar (Barnes et al., 2020), and OORD (Gadd et al., 2024) collected data under conditions such as fog, rain, and snow.
@@
The trend towards multi-sensor fusion spurred the creation of diverse and complex datasets. Datasets such as NCLT (Carlevaris-Bianco et al., 2016), M2DGR (Yin et al., 2021), NTU-VIRAL (Nguyen et al., 2022), ALITA (Yin et al., 2022), and FusionPortable (Jiao et al., 2022) also pose challenges for SLAM, given their diverse environmental appearance and structure. These datasets feature a variety of environments, including dense vegetation, open spaces, and complex buildings with multiple levels and detailed layouts. The changing lighting, seasonal foliage variations, and movement of pedestrians and vehicles add complexity to campus environments. As highlighted in NCLT dataset, these factors are crucial for life-long SLAM challenges. However, these datasets, collected via specific platforms such as unmanned ground vehicles (UGVs) and drones, fall short in showcasing diverse motion patterns, especially aggressive maneuvers.
Recent advancements in sensor technology have enabled the development of portable multi-sensor suites capable of collecting high-quality, multi-modal data in diverse environments (e.g., multi-floor buildings). Notable examples include the Newer College (Ramezani et al., 2020) and Hilti-Oxford (Zhang et al., 2022) datasets. These datasets also offer dense, high-quality 3D global maps, enabling the generation of high-rate $6$-DoF reference poses and the evaluation of mapping algorithms. The RELLIS-3D dataset (Jiang et al., 2021), though it primarily focuses on semantic scene understanding in off-road environments, also provides diverse sensor data that can be valuable for SLAM research.
2.2 Cross-Platform Datasets
@@
As SLAM research progressed from single-platform or collaborative applications, there has been a growing interest in cross-platform generalization. Building upon the specific-platform datasets discussed in Section 2.1, several datasets explore the generalization of algorithms across different platforms and scales, aiming to integrate motion characteristics from varied platforms with minimal parameter tuning for diverse scenarios. The MVSEC dataset (Zhu et al., 2018) collected multi-sensor data with diverse platforms, excluding UGV sequences. Conversely, the Nebula dataset (Reinke et al., 2022), developed during the DARPA Subterranean Challenge, includes field environments with both wheeled and legged robots, providing precise maps and trajectories. However, it lacks urban data and primarily focuses on LiDAR-based perception. The M3ED dataset (Chaney et al., 2023), although closely aligned with our objectives, lacks indoor data and platform-specific kinematic measurements, underscoring the unique contribution of our dataset.
Despite these advancements, there remains an absence of datasets that comprehensively cover a wide range of platforms, environments, and sensor modalities within a single, unified framework.
@@
Table 2: The sensors used in this dataset and their corresponding specifications. The detailed definition of ROS message type and naming of coordinate frames of each sensor are provided in the dataset website.
@@ 3 System Overview
This section presents our developed multi-sensor suite, designed for integration with various mobile platforms through plug-and-play functionality. All sensors are securely mounted on a aluminum alloy frame, facilitating a unified installation. Additionally, we detail the devices employed for collecting GT trajectories and maps.
@@ 3.1 Suite Setup and Synchronization
The Multi-Sensor Suite (MSS) integrates exteroceptive and proprioceptive sensors, including a 3D Ouster LiDAR, stereo frame and event cameras, and IMUs, as depicted in its CAD model in Fig. 1. We use two PCs that are synchronized via a Network Time Protocol (NTP) server for data collection. The primary PC processes data from the frame cameras and IMU, while the auxiliary PC handles additional data types. Both PCs are equipped with a 1TB SSD, 64GB of DDR4 memory, and an Intel i7 processor, running Ubuntu with a real-time kernel patch and employing the Robot Operating System (ROS) for data collection. This distributed architecture reduces the number of ROS nodes running on each individual PC, thereby alleviating the risk of queuing problems during data collection. After the mission, separate data collected by these two PCs are merged and postprocessed offline. The subsequent sections will elaborate on the synchronization approach and the features of each sensor.
(a) Illustration of data flow and synchronization
(b) Synchronization reduces delay in $t_{delay}$ compared with $t_{offset}$
Figure 2: (a) Illustration of data collection which shows the data flow and synchronization processes. The red arrow indicate PPS signals for synchronization, green arrows show UTC time synchronization, and blue arrows represent sensor triggering signals, and black arrows depict the flow of raw data. (b) The timing diagram for triggerable (our case) and non-triggerable sensors, illustrating the unknown time offset caused by the delay in starting data capture, the duration of data capture, and the time required for data transmission from the sensor to the PC. Our synchronization solution can reduce the time delay (i.e., $t_{offset}-t_{delay}$) but cannot address other factors, which require online time calibration algorithms.
3.1.1 Synchronization:
The synchronization process is illustrated in Fig. 2. Generally, the field-programmable gate array (FPGA) board synchronizes with the pulse-per-second (PPS) signal from the external GNSS receiver, producing higher frequency trigger signals for the IMU, stereo frame cameras, and LiDAR clock alignment. In GPS-denied environments, it utilizes its internal low drift oscillator for synchronization, achieving a time accuracy below $1ms$ between multiple trigger signals. To synchronize LiDAR and camera data, we phase-lock11endnote: 1https://static.ouster.dev/sensor-docs/image_route1/image_route2/time_sync/time-sync.html the LiDAR’s rotation so its forward-facing direction aligns with the camera’s capture timing, accounting for the continuous nature of LiDAR’s spinning data. We use the internal (Master-Slave mode) mechanism to synchronize stereo event cameras, where the left event camera is assigned as the master to send trigger signals to the right camera.
Fig. 2(b) illustrates our synchronization scheme. The FPGA sends trigger signals to connected sensors, which starts capturing data after a small delay (commonly $t_{delay}<1ms$). $t_{capture}$ is defined as the time for capturing data such as camera exposure or LiDAR rotation. It is followed by a transmission process that sends data to the host with time $t_{transmit}$. Non-triggerable sensors such as wheeled encoders use their internal clocks for timestamping, resulting in a time offset, $t_{offset}$, relative to the FPGA’s trigger signal. Our scheme cannot recover $t_{delay}$ and $t_{capture}$.